# TranDRL: A Transformer-Driven Deep Reinforcement Learning Enabled Prescriptive Maintenance Framework

### v2

@@ 10 Output recommended action $a$;
Algorithm 1 A Transformer-enabled DRL framework for maintenance selection.
@@ IV-A Case 1: Remaining Useful Life Prediction Module for a Single Machine
Transformers, rooted in natural language processing [27], have become valuable tools in time-series data analysis. This is largely due to their proficiency in understanding complex temporal sequences. Our model leverages the Transformer’s capabilities to forecast the RUL from historical time-series sensor data. The core of this approach is the self-attention mechanism, which establishes causal relations between different time steps based on query-and-key similarities, adapting to the complexity of the data, for example, self-attention can handle complex relationships in machinery sensor data when issues depend on small changes over time in predictive maintenance.
@@ IV-A1 Encoder Module
For a given time-series $\mathbf{X}=[\bm{x}_{1},\bm{x}_{2},\cdots,\bm{x}_{n}]$, where $n$ is the sequence length and $\bm{x}_{i}\in\mathbb{R}^{d_{\text{model}}}$, the Encoder Module converts $\mathbf{X}$ into hidden states $\bm{h}=[h_{1},h_{2},\cdots,h_{n}]$, with each $h_{i}\in\mathbb{R}^{d_{\text{model}}}$. Each Transformer block in this module comprises:
  • Multi-Head Self-Attention Mechanism. Given matrices for query $\mathbf{Q}\in\mathbb{R}^{n\times d_{k}}$, key $\mathbf{K}\in\mathbb{R}^{n\times d_{k}}$, and value $\mathbf{V}\in\mathbb{R}^{n\times d_{v}}$, the attention mechanism is:
    $\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{Softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d_{k}}}\right)\mathbf{V},$
    with $d_{k}$ being the dimension of queries and keys, and $d_{v}$ the dimension of values.
  • Feed-Forward Neural Network. To capture non-linear patterns, this is represented as:
@@
IV-A2 Self-Attention Mechanism
This module facilitates the model in emphasizing or downplaying certain temporal characteristics. The attention weight $\bm{\Lambda}=[\alpha]_{i,j}$ computes the weighted sum with its $(i,j)$-th element defined as:
where $\alpha_{ij}$ signifies the attention weight of the $j$-th input for the $i$-th output.
IV-A3 Decoder Module
After encoding the time-series data into hidden states, the decoder further refines this information to produce outputs which are conductive for RUL prediction. Given hidden states $\bm{h}^{\prime}$, the decoder module generates output states $\textbf{O}=\{\bm{o}_{1},\bm{o}_{2},\cdots,\bm{o}_{n}\}$. Each block in the decoder encompasses:
@@ IV-A3 Decoder Module
    where $d_{k}$ represents the dimension of the queries and keys.
  • Feed-Forward Neural Network. This module extracts patterns in the decoder states:
@@
IV-A4 Prediction Module
The prediction module utilizes the states $\bm{h}$ to estimate the RUL. This involves one or several dense layers, which can be expressed as
@@ 17             Add another residual connection followed by layer normalization;
Algorithm 2 A Transformer algorithm for a single machine.
IV-B Case 2: A Federated RUL Prediction Module for Multi-Machines
The multi-machine case is featured by a federated model, which dispatches an individual training model for similar machines. The local model takes as input the local time series data $\mathbf{X}$ and outputs an estimation of the RUL $R(t)$. By combining the strengths of multiple models extracted from different machines, the approach improves the reliability and accuracy of RUL predictions, enabling industries to make informed decisions regarding maintenance schedules and machinery operations. The averaged model for RUL prediction leverages an ensemble of predictive models to provide a more precise estimate of RUL, enhancing the efficiency and reliability of predictive maintenance in industries.
$R(t)=f_{\text{ensemble}}(\mathbf{X}),$ (4)
By integrating the predictions of multiple models, the ensemble approach refines the reliability and accuracy of the RUL predictions. This enhanced prediction quality allows industries to make precise maintenance schedules and optimize machine operations. The averaged model for RUL prediction can be represented as:
$R(t)=\frac{1}{m}\sum_{i=1}^{m}f_{i}(\mathbf{X}),$ (5)
where $m$ is the number of models in the ensemble, and $f_{i}$ represents the individual predictive model. This averaging method capitalizes on the collective strength of diverse predictive algorithms. Based on the predictive RULs, the infrastructure can conduct equipment maintenance or replacement to ensure uninterrupted and reliable functionality of the machines.
@@ IV-B Case 2: A Federated RUL Prediction Module for Multi-Machines
1 Initialize a central model with positional encoding and input embeddings; for each training round do
@@ 14       Generate output sequence $\mathbf{y}_{k}$ using the final model on local input $\mathbf{x}_{k}$;
return $\mathbf{y}_{1},\mathbf{y}_{2},\ldots,\mathbf{y}_{K}$
Algorithm 3 A Federated Transformer Algorithm for Multiple Machines
@@
The distinction between Case 1 and Case 2 lies primarily in their scale and data handling methods. Case 1 focuses on predicting the RUL for a single machine using a Transformer model, which excels in interpreting complex temporal sequences in sensor data. This approach is highly specialized for analyzing the intricate patterns of an individual machine. In contrast, Case 2 extends to multiple machines, employing a federated learning approach with an ensemble model. This method aggregates data from various machines, synthesizing diverse inputs for a broader analysis. It suits for a industry which contains multiple similar production lines or machines. Additionally, Case 2 could apply Case 1’s algorithm to each machine and then focus on multi-machine data aggregation.
IV-C Maintenance Action Recommendation Module
Rule-Based Systems. These systems are established on the basis of predetermined heuristic principles. For instance, should the RUL prediction for a component fall below a designed limit (e.g., 10 hours), an immediate maintenance notification is dispatched. On the other hand, if the RUL surpasses the limit, periodic inspections is considered sufficient.
The appropriate integration of the RUL predictions with the maintenance action recommendation module ensures that maintenance schedules is efficient and data-driven, ultimately leading to improved system reliability, longer equipment lifespan, and reduced operational costs. Thus, the predicted RUL, indicative of a system’s operational state, is assimilated into the state representation within a Markov Decision Process (MDP)-based framework. The subsequent discussions elaborate the maintenance strategy-learning procedure under the MDP framework, which includes following main components:
Environment State Space. As established, the system can be in one of $n$ distinct degradation states, with $\mathbf{S}^{t}=[S_{1}^{t},S_{2}^{t},\ldots,S_{n}^{t}]$ consisting all possible $n$ states that the system can take at time $t$.
$R(s_{i}^{t},a_{k}^{t})=\sum_{j=1}^{n}P_{i,j}^{k}R(s_{i}^{t},a_{k}^{t},s_{j}^{t+1}).$
Policy. A policy $\pi$ is a mapping from states to actions. An optimal policy $\pi^{*}$ maximizes the expected cumulative discounted reward from any starting state. The value of a state $s_{i}$ under policy $\pi$ is:
where $\gamma$ is the discount factor for future rewards.
$V^{*}(s_{i})=\max_{a\in A}\left[R(s_{i},a)+\gamma\sum_{j=1}^{n}P_{i,j}^{a}V^{*}(s_{j})\right].$
On top of this MDP-based policy derivation model, we are able to develop a DRL-based plug-and-play module for maintenance action recommendation using neural network-based function/policy approximation, by accommodating a number of state-of-the-art DRL algorithms. Algorithm 4 shows the details of leveraging different DRL algorithms, including DQN, SAC, PPO, for selecting optimal maintenance action $a$ through evaluating the performance of each policy. Details of hyperparameters setting and how to apply the algorithm to the real-world case are explained in Section V-C4.
@@ IV-C Maintenance Action Recommendation Module
Input: Exploration rate $\epsilon$ (for DQN, SAC), learning rate $\alpha$, discount factor $\gamma$ and policy update frequency (for PPO, SAC)
@@ 1 Initialize the policy network and optionally value/critic network(s) with the input hyperparameters;
3 Initialize replay buffer if applicable;
4 for each episode do
@@ 34       Periodically evaluate the performance of the current policy;
Algorithm 4 A Deep Reinforcement Learning Algorithm for Maintenance Action Recommendation
V Case Study
In this section, we perform a case study on aircraft engine maintenance using NASA’s C-MAPASS dataset [28] to evaluate our proposed TranDRL framework.
V-A Dataset
The NASA’s C-MAPSS dataset [28] contains four subsets denoted as FD001, FD002, FD003, and FD004, each having a training and test set, reflecting distinct operating and fault modes. The training data comprises inputs from three operating conditions and 21 sensors’ measurements. For training, every engine data point serves as a sample paired with its RUL. In the experiment, we applied a piecewise linear degradation model [29] for RUL determinations. The last cycle data point for each engine generally stands as the sample, with actual RUL values included in the dataset.
Figure 2: An Overview of the Temporal Fusion Transformer.
V-B Remaining Useful Life Prediction
The framework given in Algorithm 3 makes it suitable to incorporate the Transformer-based time-series event forecasting networks with other feature extraction and post-processing modules. In the case study, we adopt the Temporal Fusion Transformer (TFT) [30] in Algorithm 3 to utilizes the advantages of both recurrent neural networks and Transformers [31] into a unified encoder-decoder architecture. Compared to traditional time series forecasting models, e.g., Long Short-Term Memory networks (LSTMs) [32], TFT offers enhanced interpretability and the capability to deal with multiple prediction horizons. The structure of TFT is shown in the Figure 2 and the includes following components:
Encoder. The encoder transforms the raw time-series data into a high-dimensional space using convolutional layers and self-attention mechanisms, thereby learning rich representations of the input.
Temporal Attention. This specialized module let the model focus on different time-steps, facilitating the capture of dynamic temporal patterns and long-range dependencies.
Decoder. Tailored for multi-horizon prediction, the decoder employs an autoregressive model that recursively uses its previous forecasts as inputs for forecasting subsequent steps.
Static Inputs. Apart from time-dependent features, TFT can also integrate static inputs to provide a comprehensive context for forecasting.
Interpretable Components. To ensure transparency, TFT integrates mechanisms like gating and linear autoregressive decoders that allow for human-readable prediction interpretation.
Evaluation Results. In our comparisons, Figures 3 to 6 demonstrate the TFT algorithm’s superior performance over NBeats [33] and DeepAR [34] in predicting the RUL for a individual machine, as evidenced by consistently higher RUL values. This suggests TFT’s enhanced capability in capturing complex, time-dependent patterns in sensor data, highlighting its robustness across various machine behaviors. In contrast, Figure 7 shows that while federated learning models generate reasonably good average RUL values across multiple machines, they may not capture the intricate details as effectively as single-machine models. However, Figures 8 to 11 show a significant bias when models trained on a single machine are applied to other machines, underscoring the challenge of model generalizability in diverse operational environments. This analysis illustrates the trade-off between the granularity and specificity of single-machine models and the broader applicability but potentially lower precision of federated learning approaches, highlighting the need for context-aware modeling strategies in predictive maintenance.
Figure 3: The RUL prediction result for machine FD001.
Figure 4: The RUL prediction result for machine FD002.
Figure 5: The RUL prediction result for machine FD003.
Figure 6: The RUL prediction result for machine FD004.
Figure 7: Comparison of RUL prediction results of four machines using the averaged model.
Figure 8: Comparison of RUL prediction results of four machines using the model trained from machine FD001.
Figure 9: Comparison of RUL prediction results of four machines using the model trained from machine FD002.
Figure 10: Comparison of RUL prediction results of four machines using the model trained from machine FD003.
Figure 11: Comparison of RUL prediction results of four machines using the model trained from machine FD004.
V-C Optimal Maintenance Actions Recommendation
@@
In the RUL prediction stage, experiment results show that our proposed framework succeeds in predicting RULs using four datasets from NASA C-MAPSS dataset [28]. However, a gap still exists between the prediction results and the maintenance action decisions needed by engineers. To make action recommendation by weighing the cost and benefits, we leverage DRL algorithms as given in Algorithm 4, which helps to find the best trade-off.
V-C1 States and Actions
@@
The state space is generated from PCA features of the RUL resulted from the prediction outcomes of the NASA’s C-MAPSS Aircraft Engine Simulator dataset. These PCA features provide a compact and important part of the machine’s current health condition. The action space contains three actions:
@@
V-C2 Probabilities Model
In our study, the transition and action cost matrices are empirically calibrated using the NASA C-MAPSS dataset. Specifically, each entry in the state transition matrix $p(S_{t+1}|S_{t},a_{t})$ reflects the probability of transitioning from state $S_{t}$ to $S_{t+1}$ when action $a_{t}$ is taken, and these probabilities are derived through statistical analyses of state sequences and their corresponding RUL values in the C-MAPSS dataset. Moreover, action-related costs, such as maintenance expenses or energy consumption, are grounded in real-world data from this dataset. In this MDP framework, the transition matrix has a dimension $(10,3,10)$, representing transitions between $10$ discretized states given $3$ actions, all closely reflecting scenarios that commercial aero-engines might encounter.
V-C3 Reward Model
The objective of maintenance actions recommend algorithms is not only to maximize the RUL but also to minimize the operational costs and downtime. To fulfill these multi-objective requirements, we propose a reward function that combines these elements in a weighted manner. The reward function $R(S_{t},a_{t},S_{t+1})$ is formulated as:
@@ V-C3 Reward Model
  • $\text{Cost}(a_{t})$ denotes the operational or maintenance cost incurred by taking action $a_{t}$. This cost is modeled based on experiences.
@@
V-C4 Hyperparamters
For each DRL algorithm embedded in Algorithm 4, we carefully choose its corresponding hyperparameters based on empirical experiments. Details for each DRL algorithm are given as follows.
Deep Q-network (DQN). DQN reduces the computational burden of tabular Q-learning by employing a deep neural network to approximate the state-action value function [35]. It also introduces experience replay to improve learning stability and efficiency. The learning rate is set to $10^{-4}$, the discount factor is set to $0.99$, and the replay buffer size is set to $10^{6}$ experiences. We use a batch size of $512$ and update the target network every $50$ steps. The exploration-exploitation trade-off is controlled by an $\epsilon$ value that started at $1.0$ and decayed over time to a final value of $0.01$. Additionally, we use a neural network with two hidden layers of $256$ units each, and we utilize the Adam optimizer with a learning rate of $10^{-4}$.
Proximal Policy Optimization (PPO). PPO ensures more stable and efficient policy updates by employing a clipped objective function, thereby avoiding large, destabilizing updates to the policy [36]. We set the learning rate for the actor and critic networks to $0.0001$ and the discount factor to $0.99$. A clip range of $0.2$ is used to limit the updates to the policy, and we apply a value function coefficient of $0.5$ to balance the importance of the policy and the value function. Additionally, we set the entropy coefficient to $0.01$ to encourage exploration and a batch size as $64$. Then, we optimize the policy every 10 epochs and use the Adam optimizer with a learning rate of $0.0001$. Finally, GAE parameter is set to $0.95$ to estimate the advantage function.
Soft Actor-Critic (SAC). SAC is particularly effective for problems with high-dimensional action spaces [37]. SAC aims to maximize an objective function that balances both the expected return and the entropy of the policy, making it more robust and efficient. The learning rate for the actor and critic networks is set to $0.0003$, and we use a discount factor of $0.99$. Then, we apply a temperature parameter alpha of $0.2$ to balance the entropy term in the objective function. The target entropy is set to $-2.0$, and we use a replay buffer of $10^{6}$ experiences with a batch size of $256$. We also apply a soft update of the target networks with a rate of $0.005$, and we initialize the log standard deviation of the policy with a value of $-2$. Finally, we train the algorithm for $10^{5}$ time steps and use the Adam optimizer with a learning rate of $0.0001$.
V-D Results and Discussion
Figure 12: The rewards for $10^{5}$ episodes during the training of DQN, PPO, and SAC algorithms in our designed environment.
Figure 13: The rewards for $10^{6}$ episodes during the training of DQN, PPO, and SAC algorithms in our designed environment.
We compare the performance of DQN, PPO, and SAC on our designed task in Figures 1213. The evaluation metric is the total rewards obtained by the agent during a fixed number of episodes. The experimental results show that PPO outperforms both DQN and SAC, achieving higher total rewards on the task. The average total rewards obtained by PPO are $55,000$, while DQN and SAC achieve average total rewards of $50,000$ and $45,000$, respectively. However, DQN converges faster than PPO and SAC.
In general, PPO and SAC are considered to be more sample-efficient than DQN, which means that they can achieve comparable performance with fewer samples. This is because PPO and SAC use more efficient updates that make better use of the available data. In contrast, DQN can be more sensitive to the choice of hyperparameters and the amount of data used during training, which can affect its convergence speed. However, the performance of these algorithms can vary depending on the specific task and the quality of the implementation. In some cases, DQN may converge faster than PPO or SAC, especially if the action space is discrete and the observations are low-dimensional. On the other hand, if the action space is continuous or the observations are high-dimensional, PPO and SAC may be more suitable and converge faster.
Overall, these experimental results demonstrate the importance of selecting the appropriate DRL algorithm and hyperparameters for a specific task to achieve the optimal performance.
Figure 14: Differences between RL and RLHF.
V-E Reinforcement Learning with Human Feedback
Incorporating human expertise into the reinforcement learning process helps to resolve some of the challenges encountered with traditional RL methods. Especially, in scenarios where real-world rewards are hard to quantify or can be misleading, human feedback provides a way to guide the learning process more effectively. The comparison between RL and RLHF is presented in Fig. 14.

### v3

@@ 12 Output recommended action $a$;
Algorithm 1 TranDRL:A Transformer-enabled DRL framework for optimal maintenance action recommendation.
@@ IV-A Case 1: Remaining Useful Life Prediction Module for a Single Machine
Transformers, originating from natural language processing [25], have become valuable tools in time-series data analysis, due to their proficiency in understanding complex temporal sequences. Our proposed Algorithm 2 leverages a transformer to forecast the RUL from historical time-series sensor data. The core of this approach is the self-attention mechanism, which establishes causal relations between different time steps based on query-and-key similarities, adapting to the complexity of the data, for example, self-attention can handle complex relationships in sensor data while issues depend on small changes over time in predictive maintenance.
@@ IV-A1 Encoder Module
For a given time-series $\mathbf{X}=[\bm{x}_{1},\bm{x}_{2},\cdots,\bm{x}_{n}]$, where $n$ is the sequence length and $\bm{x}_{i}\in\mathbb{R}^{d_{\text{model}}}$, the Encoder module converts $\mathbf{X}$ into hidden states $\bm{h}=[h_{1},h_{2},\cdots,h_{n}]$, with each $h_{i}\in\mathbb{R}^{d_{\text{model}}}$. This module comprises an input embedding layer and a number of $L$ identical self-attention blocks as follows:
  • Input Embedding. The input data in $\mathbf{X}$ are processed with a learned embedding layer and converted to vectors of dimension $d_{\text{model}}$. The embedding results are then added with a learned positional encoding of the same dimension to reflect the temporal difference between the input samples.
  • Multi-head Self-attention Mechanism. Given the matrices for query $\mathbf{Q}\in\mathbb{R}^{n\times d_{k}}$, key $\mathbf{K}\in\mathbb{R}^{n\times d_{k}}$, and value $\mathbf{V}\in\mathbb{R}^{n\times d_{v}}$, respectively, the attention score is computed as follows:
    $\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{Softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d_{k}}}\right)\mathbf{V},$ (1)
[…]
    since $\alpha_{ij}$ signifies the attention weight of the $j$-th input for the $i$-th output. With multiple heads, the queries, keys, and values are learned in parallel linear projections, and then concatenated as the output of this layer.
  • Feed-Forward Neural Network (FFN). A single-layer FFN is utilized to capture non-linear patterns. This layer is represented as:
@@
IV-A2 Decoder Module
After encoding the time-series data into hidden states, the decoder further refines this information to produce outputs that are conducive to RUL prediction. Given hidden states $\bm{h}^{\prime}$, the decoder module generates output states $\textbf{O}=\{\bm{o}_{1},\bm{o}_{2},\cdots,\bm{o}_{n}\}$. Similar to the encoder, the decoder module is also composed of a number of $L$ identical blocks. Each block in the decoder encompasses:
  • Masked Multi-head Self-attention Mechanism. This layer takes the output embedding as the input and is the same as the multi-head self-attention mechanism in the encoder module. Masking is applied to make sure that the newly generated output only depends on the historical outputs.
@@ IV-A2 Decoder Module
    where $d_{k}$ represents the dimension of the queries and keys. Being different from (1), the values of $\mathbf{K}_{e}$ and $\mathbf{V}_{e}$ are derived from the output of the encoder and $\mathbf{Q}_{d}$ is derived from the output embedding of the decoder or the output of the previous block in the decoder.
  • Feed-forward Neural Network. This module extracts patterns in the decoder states and takes the same form as (3) in the encoder:
@@
We note that each attention layer or FFN layer in the encoder and the decoder is accompanied with a residual connection and their output is processed with layer normalization (see Algorithm 2).
IV-A3 Prediction Module
The prediction module utilizes the states $\bm{h}$ to estimate the RUL. This involves one or several dense FFN layers, each layer of which can be expressed as
@@ 17             Add another residual connection followed by layer normalization;
Algorithm 2 A RUL prediction algorithm for a single machine
@@
1
2Initialize a central model with positional encoding and input embeddings;
3for each training round do
@@ 18       Generate output sequence $\mathbf{y}_{k}$ using the final model on local input $\mathbf{x}_{k}$;
21
return $\mathbf{y}_{1},\mathbf{y}_{2},\ldots,\mathbf{y}_{K}$.
Algorithm 3 A federated RUL prediction algorithm for multiple machines
IV-B Case 2: A Federated RUL Prediction Module for Multi-Machines
The multi-machine case is featured by a federated model, which dispatches an individual training model for $K$ similar machines. A local model $k$ takes as input the local time series data $\mathbf{X}_{k}$ and outputs an estimation of the RUL $R_{k}(t)$ regarding the machine it concerns. With sensors deployed in the same way, different local models observe time-series inputs of the same dimension and share the same prediction label set. By combining the strengths of multiple models extracted from different machines, the approach improves the reliability and accuracy of RUL predictions, enabling industries to make informed decisions regarding maintenance schedules and machinery operations. The global model for RUL prediction leverages an ensemble of predictive models to provide a more precise estimate of RUL, enhancing the efficiency and reliability of predictive maintenance in industries.
$R(t)=f_{\text{ensemble}}(\mathbf{X}),$ (5)
By integrating the predictions of multiple models, the ensemble approach refines the reliability and accuracy of the RUL predictions. This enhanced prediction quality allows industries to make precise maintenance schedules and optimize machine operations. The global model for RUL prediction can be represented as
$R(t)=\frac{1}{K}\sum_{k=1}^{K}f_{k}(\mathbf{X}),$ (6)
where $K$ is the number of models in the ensemble, and $f_{k}$ represents the individual predictive model. This averaging method capitalizes on the collective strength of diverse predictive algorithms. Based on the predictive RULs, the infrastructure can conduct equipment maintenance or replacement to ensure uninterrupted and reliable functionality of the machines.
@@ IV-B Case 2: A Federated RUL Prediction Module for Multi-Machines
The distinction between Case 1 and Case 2 lies primarily in their scale and data handling methods. Case 1 focuses on predicting the RUL for a single machine using a Transformer model, which excels in interpreting complex temporal sequences in sensor data. This approach is highly specialized for analyzing the intricate patterns of an individual machine. In contrast, Case 2 extends to multiple machines, employing a federated learning approach with an ensemble model. This method aggregates data from various machines, synthesizing diverse inputs for a broader analysis. It suits an industry that contains multiple similar production lines or machines. Additionally, Case 2 can suitably apply Case 1’s algorithm to each machine and then focus on multi-machine data aggregation.
@@
Input: Exploration rate $\epsilon$ (for DQN, SAC), learning rate $\alpha$, discount factor $\gamma$ and policy update frequency (for PPO, SAC).
@@ 2 Initialize the policy network and optionally value/critic network(s) with the input hyperparameters;
4 Initialize experience replay buffer if applicable;
5
6for each episode do
@@ 38       Periodically evaluate the performance of the current policy;
Algorithm 4 A DRL-based algorithm for optimal maintenance action recommendation
IV-C Deep Reinforcement Learning for Maintenance Action Recommendation
Rule-Based Systems. These systems are established on the basis of predetermined heuristic principles. For instance, should the RUL prediction for a component fall below a designed limit (e.g., 10 hours), an immediate maintenance notification is dispatched. On the other hand, if the RUL surpasses the limit, periodic inspections are considered sufficient.
The appropriate integration of the RUL predictions with the maintenance action recommendation module ensures that maintenance schedules are efficient and data-driven, ultimately leading to improved system reliability, longer equipment lifespan, and reduced operational costs. Thus, the predicted RUL, indicative of a system’s operational state, is assimilated into the state representation within a Markov Decision Process (MDP)-based framework. The subsequent discussions elaborate on the maintenance strategy-learning procedure under the MDP framework, which includes the following components:
Environment State Space. As established, the system can be in one of $n$ distinct degradation states, with $\mathbf{S}^{t}=\{S_{1}^{t},S_{2}^{t},\ldots,S_{n}^{t}\}$ consisting all possible $n$ states that the system can take at time $t$.
$R(s_{i}^{t},a_{k}^{t})=\sum_{j=1}^{n}P_{i,j}^{k}R(s_{i}^{t},a_{k}^{t},s_{j}^{t+1}).$
Policy. A policy $\pi$ is a (probabilistic) mapping from states to actions. An optimal policy $\pi^{*}$ maximizes the expected cumulative discounted reward from any starting state. The value of a state $s_{i}$ under policy $\pi$ is:
where $\gamma$ is the discount factor for future rewards.
$V^{*}(s_{i})=\max_{a\in A}\left[R(s_{i},a)+\gamma\sum_{j=1}^{n}P_{i,j}^{a}V^{*}(s_{j})\right].$ (7)
On top of this MDP-based policy derivation model, we can develop a DRL-based plug-and-play module for maintenance action recommendation using neural network-based function/policy approximation, by accommodating many state-of-the-art DRL algorithms. Algorithm 4 shows details of leveraging different DRL algorithms, including DQN, SAC, and PPO (see our later discussion for the details of each algorithm), for selecting optimal maintenance action $a$ through evaluating the performance of each policy.
@@ IV-C Deep Reinforcement Learning for Maintenance Action Recommendation
In the RUL prediction stage, experiment results in Section V show that our proposed framework succeeds in predicting RULs using four sub-datasets in the NASA C-MAPSS open dataset [26]. However, a gap still exists between the prediction results and the maintenance action decisions needed by engineers. To make action recommendations by weighing the cost and benefits, we leverage DRL algorithms as given in Algorithm 4, which helps to find the best trade-off.
IV-C1 States and Actions
@@
The state space is generated from PCA features of the RUL resulting from the prediction outcomes of NASA’s C-MAPSS Aircraft Engine Simulator dataset. These PCA features provide a compact and important part of the current health condition of the machine. The action space contains three actions:
@@
IV-C2 Probabilities Model
For the purpose of evaluating the performance of our proposed algorithm, the transition and action cost matrices are empirically calibrated using the NASA C-MAPSS dataset to generate the controllable ground truth. Specifically, each entry in the state transition matrix $p(s_{t+1}|s_{t},a_{t})$ reflects the probability of transitioning from state $s_{t}$ to $s_{t+1}$ when action $a_{t}$ is taken, and these probabilities are derived through statistical analyses of state sequences and their corresponding RUL values in the C-MAPSS dataset. Moreover, action-related costs, such as maintenance expenses or energy consumption, are grounded in real-world data from this dataset. In this MDP framework, the transition matrix has a dimension $(10,3,10)$, representing transitions between $10$ discretized states given $3$ actions, all closely reflecting scenarios that commercial aero-engines might encounter.
IV-C3 Reward Model
The objective of maintenance actions recommends algorithms not only to maximize the RUL but also to minimize operational costs and downtime. To fulfill these multi-objective requirements, we propose a reward function that combines these elements in a weighted manner. The reward function $R(s_{t},a_{t},s_{t+1})$ is formulated as
@@ IV-C3 Reward Model
  • $\text{Cost}(a_{t})$ denotes the operational or maintenance cost incurred by taking action $a_{t}$. This cost is modeled based on experience.
@@
IV-D Integrate with Human Feedback for Maintenance Action Recommendation